# That is Unacceptable: the Moral Foundations of Canceling

### v1

@@ 5.2 STUDY 2: Do Different LLMs Align with Different Moral Profiles in Evaluating Canceling Attitudes?
Figure 2: LLM variability in alignment on the unacceptability task.
@@
From an in-depth observation of Figure 2 both common and specific patterns among models emerge. LLMs show different degrees of alignment with specific moral clusters. OLMo classifies unacceptability with a highest F-score against annotations provided by moral cluster_0 in $5$ cases out of $6$ while OPT-IML shows an opposite alignment, obtaining a highest F-score against moral cluster_1 annotations in $4$ cases out of $5$. The alignment of LLMs with annotators broken down by gender appears to systematically lean towards men. BLOOMZ and DeepSeek always obtain the highest F-1 scores against men’s annotations, OPT-IML in $5$ cases out of $6$. Models that align the most with women are OLMo and Ministral, which both score a highest F-score against them in $3$ cases out of $6$. The analysis of stakeholders shows a general higher alignment of LLMs with students, which are the most represented group in our pool of annotators. In particular, BLOOMZ always aligns with this category of stakeholders. On the opposite, activists are the stakeholders against which models are more misaligned. Specifically, Llama and Ministral obtain the lowest F-score against annotations provided by this group in $5$ cases out of $6$.
@@ 6 Discussion
This event-based variation explains human disagreement in the evaluation of canceling attitudes. Rather than operating at a general level, morality, gender, and stakeholder type drive the disagreement of annotators against specific celebrities. In this context, morality appears to have a significant influence, especially in the evaluation of unacceptability. In $4$ cases out of $6$ annotators with different moral profiles show a statistically significant variation (Table 2) in judging unacceptability. This highlights the importance of developing resources that better represent the communicative context in which potentially harmful content is spread.
@@ 7 Conclusion
In this paper, we presented the Canceling Attitude DEtection (CADE) dataset, a corpus specifically designed to investigate how people with different moral views of the world evaluate this phenomenon. The corpus, which includes six canceling incidents gathered from YouTube, has been annotated by annotators belonging to three categories of relevant stakeholders for this social issue (activists, researchers, and students) whose moral profiles have been obtained through the MFT questionnaires.
Future work will be devoted to expanding the study of canceling attitudes by including in the analysis different categories of targets (e.g., organizations). Additionally, we will adopt an intersectional approach to better understand how morality intersects with other sociodemographic factors in determining annotators’ disagreement, and to which extent LLMs are able to grasp this complex issue.
@@
While this work moves towards the development of participatory approaches to Natural Language Processing, the annotator pool is not balanced, especially leaning towards the White and Educated population. Moreover, although the chosen events had worldwide coverage, the majority of the annotators do not come from the same social background as the target celebrities. This made it possible to carry out the experiment with people from different nationalities, who, however, experienced cancel culture more as spectators than actors. In the future, we plan to expand the annotation lab to a more diverse group of annotators along all the sociodemographic axes. Finally, the current design involved watching only one video per celebrity, with exposure to the framing implicit in the political orientation of the chosen news source. While we aimed to ensure political diversity across the six selected videos, in the future, we intend to achieve this within the same canceling event.
@@
This research relies on the voluntary work of those who participated in the Annotation Labs. All the involved annotators freely accepted to take part to the laboratory, for which no compensation was provided. We adopted all the measures to protect data privacy and safeguard personal information. The work has been approved by the Ethics Committee of the institution of one of the authors.
@@ Appendix A Annotator’s demographics
Table 3 shows the demographic information about the annotators. We additionally asked them to indicate their nationality and first language. 38 are Italian, 2 are Spanish, 2 are Chinese, and there is one person for each of the following nationalities: Italian-Argentinian, Italian-Romanian, Iranian, Greek, Russian, Kazakhstan, Indian, Moldovan, Persian, Dutch, Romanian. As regards their mother tongue: 38 chose Italian, 3 Spanish, Russian, Chinese, 2 Persian, Romanian and Greek, and 1 Hebrew, Hindi, Farsi and Dutch.
@@ Appendix B Instructions for the annotation process
Figure 3 shows the instructions provided to the annotators.
@@ Appendix C Annotation materials
Table 4 presents all the information about the selected YouTube videos.
@@
Table 5 reports the detailed sample composition and Inter Annotator Agreement.
Samples #Researchers #Students #Activists #Texts #Annotations $\alpha$ Stance $\alpha$ Acceptability
@@ Appendix D Checkout Questionnaire
This questionnaire was not mandatory and received $33$ answers.
Appendix E LLM implementation details

### v3

@@ 5.2 STUDY 2: Do Different LLMs Align with Different Moral Profiles in Evaluating Canceling Attitudes?
Figure 2: LLM variability in alignment on the unacceptability task. The color bar represents F1-scores, with darker shades indicating higher values
@@
From an in-depth observation of Figure˜2 both common and specific patterns among models emerge. LLMs show different degrees of alignment with specific moral clusters. OLMo classifies unacceptability with a highest F-score against annotations provided by moral cluster_0 in $5$ cases out of $6$ while OPT-IML shows an opposite alignment, obtaining a highest F-score against moral cluster_1 annotations in $4$ cases out of $5$. The alignment of LLMs with annotators broken down by gender appears to systematically lean towards men. BLOOMZ and DeepSeek always obtain the highest F-1 scores against men’s annotations, OPT-IML in $5$ cases out of $6$. Models that align the most with women are OLMo and Ministral, which both score a highest F-score against them in $3$ cases out of $6$. The analysis of stakeholders shows a general higher alignment of LLMs with students, which are the most represented group in our pool of annotators. In particular, BLOOMZ always aligns with this category of stakeholders. On the opposite, activists are the stakeholders against which models are more misaligned. Specifically, Llama and Ministral obtain the lowest F-score against annotations provided by this group in $5$ cases out of $6$.
@@ 6 Discussion
This event-based variation explains human disagreement in the evaluation of canceling attitudes. Rather than operating at a general level, morality, gender, and stakeholder type drive the disagreement of annotators against specific celebrities. In this context, morality appears to have a significant influence, especially in the evaluation of unacceptability. In $4$ cases out of $6$, annotators with different moral profiles show a statistically significant variation (Table˜2) in judging unacceptability. This highlights the importance of developing resources that better represent the communicative context in which potentially harmful content is spread.
@@ 7 Conclusion
In this paper, we presented the Canceling Attitude DEtection (CADE) dataset, a corpus specifically designed to investigate how people with different moral views of the world evaluate this phenomenon. The corpus, which includes six canceling incidents gathered from YouTube, has been annotated by annotators belonging to three categories of relevant stakeholders for this social issue (activists, researchers, and students) whose moral profiles have been obtained through the MFT questionnaire.
Future work will be devoted to expanding the study of canceling attitudes by including different categories of targets (e.g., organizations), as well as an extensive analysis of how unacceptability varies in respect to the stance, together with used linguistic strategies. Additionally, we will adopt an intersectional approach to better understand how morality intersects with other sociodemographic factors in determining annotators’ disagreement, and to which extent LLMs are able to grasp this complex issue. To this end, we plan to analyse differences between models of the same family with a varying number of parameters, conducting extensive research on LLMs alignment.
@@
While this work moves towards the development of participatory approaches to Natural Language Processing, the annotator pool is not balanced, especially leaning towards the White and Educated population. Moreover, although the chosen events had worldwide coverage, the majority of the annotators do not come from the same social background as the target celebrities. This made it possible to carry out the experiment with people from different nationalities, who, however, experienced cancel culture more as spectators than actors. In the future, we plan to expand the annotation lab to a more diverse group of annotators along all the sociodemographic axes. In acknowledging the low agreement resulting from the annotation in Section Section˜4.2, we plan to compare these results with other disaggregated corpora, adopting metrics that emphasize the emergence of disagreement factors (Passonneau and Carpenter, 2014). Finally, the current design involved watching only one video per celebrity, with exposure to the framing implicit in the political orientation of the chosen news source. While we aimed to ensure political diversity across the six selected videos, in the future, we intend to achieve this within the same canceling event.
@@
This research relies on the voluntary work of those who participated in the Annotation Labs. All the involved annotators freely accepted to take part to the laboratory, for which no compensation was provided. We adopted all the measures to protect data privacy and safeguard personal information of both YouTube users whose comments were collected, and annotators who participated to the task. The work has been approved by the Ethics Committee of the institution of one of the authors.
We acknowledge that sharing the research objectives in advance can be a source of bias. However, we decided to prioritise the annotators’ awareness about the technology they were building, providing an overview of the work. We made an effort to ensure the direct participation of those who contributed to the construction of the dataset, not only through annotations but also with their feedback, trying to limit the negative aspects of this choice.
@@
Author Contribution Statement
  • Project conception: Soda Marem Lo, Marco Antonio Stranisci.
  • Lead: Soda Marem Lo.
  • Supervision: Marco Antonio Stranisci.
  • Experiments with LLMs: Oscar Araque.
  • Research design and paper writing: Oscar Araque, Soda Marem Lo, Rajesh Sharma, Marco Antonio Stranisci.
Acknowledgements
Oscar Araque would like to acknowledge the funding of funded by the Spanish Ministry for Economic Affairs and Digital Transformation and by the European Union - NextGenerationEU within the Programme UNICO I+D Cloud (TSI-063100-2022-0002); as well as the funding of the project CPP2023-010437 financed by the MCIN / AEI / 10.13039/501100011033 / FEDER, UE. This work has been supported by the Madrid Government (Comunidad de Madrid-Spain) under the Multiannual Agreement 2023-2026 with Universidad Politécnica de Madrid in the Line A, Emerging PhD researchers (project MORA, DOCTORES-EMERGENTES-24-9UMLXZ-37-IIGW).
The work of Soda Marem Lo was partially supported by “HARMONIA” project - M4-C2, I1.3 Partenariati Estesi - Cascade Call - FAIR - CUP C63C22000770006 - PE PE0000013 under the NextGenerationEU programme and by aequa-tech.
We thanks students of Language Technologies and Digital Humanities and Hate Trackers activists for their participation in the annotation lab.
@@ Appendix A Annotator’s demographics
Table˜4 shows the demographic information about the annotators. We additionally asked them to indicate their nationality and first language. 38 are Italian, 2 are Spanish, 2 are Chinese, and there is one person for each of the following nationalities: Italian-Argentinian, Italian-Romanian, Iranian, Greek, Russian, Kazakhstan, Indian, Moldovan, Persian, Dutch, Romanian. As regards their mother tongue: 38 chose Italian, 3 Spanish, Russian, Chinese, 2 Persian, Romanian and Greek, and 1 Hebrew, Hindi, Farsi and Dutch.
@@ Appendix B Instructions for the annotation process
Figure˜3 shows the instructions provided to the annotators.
@@ Appendix C Annotation materials
Table˜5 presents all the information about the selected YouTube videos.
@@
Table˜6 reports the detailed sample composition and Inter Annotator Agreement.
Samples #Researchers #Students #Activists #Texts #Annotations $\alpha$ Stance $\alpha$ Unaceptability
@@ Appendix D Checkout Questionnaire
This questionnaire was not mandatory and received $33$ answers. Overall, the majority of annotators found that the annotation topic was related to the Moral Foundations Theory and eight of them wrote that they have been influenced by filling the MFT questionnaire before the annotation task. In a future iteration of this work we will administer the questionnaire after the annotation. A second relevant aspect emerging from the checkout survey is that some comments are not easy to be interpreted for non-native English speakers since the context in which the controversial events had place is strongly US-centric. We will integrate the existing corpus with annotations focused on incidents from different cultural contexts. Results of the checkout questionnaire are available in the project folder [^5] https://github.com/aequa-tech/canceling-attitudes/blob/main/questionnaires/checkout_questionnaire.csv
Appendix E Corpus statistics
The stance of the comment was labeled as Neutral $2,626$ times, Attack $4,363$, and Defend $4,821$ times. Regarding the acceptability, going on a scale from 1 (totally acceptable) to 4 (totally unacceptable), the labels were assigned as follows: [1] Polite argumentation $4,539$ times, [2] (Heated) criticism $3,532$ times, [3] Attack against a specific target $1,990$ times, [4] Trashing and bullying $1,477$ times.
Appendix F Example of annotated comments
Table˜7 shows a comment for each possible combination of the two annotation dimensions (stance and unacceptability).
Stance Unacc. Celebrity Text
Attack 1 Ellen DeGeneres So many things that go behind the scenes
2 Lizzo Ain’t that the pot calling the kettle black
3 Ellen DeGeneres Cancel her
4 Ellen DeGeneres I never liked Ellen, she is a stupid overhyped KAREN
Defend 1 Andrew Tate Maybe AI voice?
2 Andrew Tate The Matrix is after him—“The Matrix” being others jealous of him.
3 Kanye West Speaking out against the Synagogue of Satan will never end good
4 J.K. Rowling 85% of the world agrees with JK Rowling. The other 15% have mental health issues.
Neutral 1 Ellen DeGeneres Bro, I wasn’t even born in 2003. The show is older than me
2 Halle Bailey Race swapping is racist.
3 Lizzo They have you so duped!! Strangely at the same time they posted this all other stations were posting the the new Trump court orders. They steer you and control you.
4 Lizzo Is my mind playing tricks on me or is the dude’s head in the pin-stripe suit kind of big for his body?
Table 7: Examples of annotated texts across all combinations of stance and unacceptability. Labels reflect full or majority agreement among annotators.
Appendix G LLM implementation details